On Mon Sept 23 (for the usual T+1 settlement), a T-bill maturing Oct 24 is quoted at a price of 99.85% of face. What is the bank discount yield?
Answer #1

Let's assume the face value of t-bill is $10,000. so price is $10,000*99.85% = $9,985.

bank discount yield = [(Face value - Price) / Face value] x 360/maturity

between Sept 23 and Oct 24, there are 31 days.

bank discount yield = [($10,000 - $9,985)/$10,000)*360/31 = ($15/$10,000)*360/31 = 0.0015*360/31 = 0.54/31 = 0.0174 or 1.74%

bank discount yield is 1.74%.